1. Use Raoult's law to calculate the vapor pressure of a mixture of hexane (C6H14, P=32.0 torr) and heptane (C7H14, P=18.0 torr) in which the mole fraction fo heptane is 0.80.

a. 32.0 torr b. 11.0 torr c. 20.8 torr d. 18.0 torr

2. Biphenyl is a symmetrical polyaromatic hydrocarbon with the molecular formula C12H10. How many fused aromatic rings does biphenyl have in its molecular structure?

a. 1 b. 2 c. 3 d. none

3. A molecule with the R-(C=O)-O-C-R bonding arrangement is called an ___.

a. alcohol b. ester c. amine d. ether
Answer #1

1.

Let pT = Total vapor pressure of the solution,

According to Raoult's law:

pT = phexaneXhexane + pheptaneXheptane

Where:

phexane and pheptane are partial vapor pressures of hexane and heptane, respectively.

Xhexane and Xheptane are the mole fractions of hexane and heptane in the liquid phase, respectively.

Given, phexane = 32 torr, pheptane = 18 torr, Xhexane = 0.20 and Xheptane = 0.80, Xhexane = 1 - Xheptane = 0.20 torr

pT = 32 x 0.20 + 18 x 0.80 = 6.4 + 14.4 =   20.8 torr.

Answer is C         

2. Biphenyl structure is It has zero fused aromatic rings. Fused aromatic rings should share one or more than one C-C bond of